OpenStudy (anonymous):

Angelo took a group of people to the movies and spent $80 on tickets. Adult tickets cost $10 and child tickets cost $6. The equation to represent his cost is 10A + 6C = $80. If there were 5 children in the group, how many adult tickets did he purchase? Show your work and use a complete sentence to explain your solution in the context of the original problem.

OpenStudy (anonymous):

we don't know A 10A + 6C = 80 , number of children = 5 10A + 6*5 = 80 solve the equation for number of adults

OpenStudy (anonymous):

10A + 6*5 =80 10A +30=80 -30 -30 10A = 50 /10 /10 A=5 So... 5 adults? @JOELMATHEWS1234

OpenStudy (anonymous):

YEP @Jimboslice

OpenStudy (anonymous):

Thank you !

OpenStudy (anonymous):

You're welcome @@Jimboslice